# Persistent Homology for High-dimensional Data Based on Spectral Methods

### v2

@@ 1 Introduction
Algebraic topology can describe the shape of a continuous manifold. In particular, it can detect if a manifold has holes, using its so-called homology groups [33]. For example, a cup has a single one-dimensional hole, or ‘loop’ (its handle), whereas a football has a single two-dimensional hole, or ‘void’ (its hollow interior). These global topological properties are often helpful for understanding an object’s overall structure. However, real-world datasets are typically given as point clouds, a discrete set of points sampled from an underlying manifold. In this setting, true homologies are trivial, as there is one connected component per point and no holes whatsoever; instead, persistent homology can be used to find holes in point clouds and to assign an importance score called persistence to each [22, 77]. Holes with high persistence are indicative of holes in the underlying manifold. Persistence homology has been successfully applied in machine learning pipelines, for instance for gait recognition [41], instance segmentation [38], and protein binding [73], as well as for neural network analysis [58].
Persistent homology works well for low-dimensional data [67] but we find that it has difficulties in high dimensionality. If data points are sampled from a low-dimensional manifold embedded in a high-dimensional ambient space (‘manifold hypothesis’), then the measurement noise typically affects all ambient dimensions. In this setting, traditional persistent homology is not robust against even low levels of noise. On a dataset as simple as a circle in $\mathbb{R}^{50}$, persistent homology based on the Euclidean distance between noisy points can fail to find the correct loop (Figure 1). The aim of our work is to find alternatives to traditional persistent homology that can robustly detect the correct topology despite high-dimensional noise.
@@
Our contributions are: [^1] Our code is available at https://github.com/berenslab/eff-ph/tree/arxiv-v2.
@@
Figure 1: a. 2D PCA of a noisy circle ($\sigma=0.25$, radius 1) in $\mathbb{R}^{50}$. Overlaid are representative cycles of the most persistent loops. b. Persistence diagrams using Euclidean distance and the effective resistance. c. Loop detection scores of persistent homology using effective resistance and Euclidean distance. d, e. UMAP and $t$-SNE reproduced the loop structure in 2D.
@@ 2 Related work
Persistent homology has long been known to be sensitive to outliers [14] and several extensions have been proposed to make it more robust. The main idea of most of these suggestions is to replace the Euclidean distance with a different distance matrix, before running persistent homology. Bendich et al. [6] suggested to use diffusion distances [18], but their empirical validation was limited to a single dataset in 2D. Anai et al. [2] suggested to use the distance-to-measure (DTM) [14] and Fernández et al. [23] proposed to use Fermat distances [29]. Vishwanath et al. [69] introduced persistent homology based on robust kernel density estimation, an approach that itself becomes challenging in high dimensionality. All of these works focused on low-dimensional datasets ($<$10D, mostly 2D or 3D), while our work specifically addresses the challenges of persistent homology in high dimensions. Our work aligns with the concurrent, more formal approach of Hiraoka et al. [36].
Below, we will recommend using effective resistance and diffusion distances for persistent homology in high-dimensional spaces. Both of these distances, as well as the shortest path distance, have been used in combination with persistent homology to analyze the topology of graph data [54, 32, 1, 66, 10, 47]. Shortest paths on the $k$NN graph were also used by Naitzat et al. [52] and Fernández et al. [23]. Motivated by the performance of UMAP [45] for dimensionality reduction,  Gardner et al. [26] and Hermansen et al. [34] used UMAP affinities to define distances for persistent homology.
Effective resistance is a well-established graph distance [21, 25]. A correction, more appropriate for large graphs, was suggested by von Luxburg et al. 2010a, 2014. When speaking of effective resistance, we mean this corrected version, if not otherwise stated. It has not yet been combined with persistent homology. Conceptually similar diffusion distances [18] have been used in single-cell RNA-sequencing data analysis, for dimensionality reduction [48], trajectory inference [30], feature extraction [15], and hierarchical clustering, similar to 0D persistent homology [8, 40].
@@
Figure 2: a. Persistent homology applied to a noisy circle ($n=10$) in 2D. Dotted lines show the edges that lead to the birth / death of two loops (Section 3). b. The corresponding persistence diagram with two detected 1D holes (loops). Our hole detection score measures the gap in persistence between the first and the second detected holes (Section 7).
@@ 3 Background: persistent homology
Persistent homology consists of a set of holes for each dimension. We limit ourselves to loops and voids. Each hole has associated birth and death times $(\tau_{b},\tau_{d})$, i.e., the first and last filtration value $\tau$ at which that hole exists. Their difference $p=\tau_{d}-\tau_{b}$ is called the persistence of the hole and quantifies its prominence. The birth and death times can be visualized as a scatter plot (Figure 2b), known as the persistence diagram. Points far from the diagonal have high persistence. This process is illustrated in Figure 2 for a noisy sample of $n=10$ points from a circle $S^{1}\subset\mathbb{R}^{2}$. At $\tau_{1}$ a small spurious loop is formed thanks to the inclusion of the dotted edge, but it dies soon afterwards. The ground-truth loop is formed at $\tau_{2}$ and dies at $\tau_{3}$, once the hole is completely filled in by triangles. Both loops (one-dimensional holes) found in this dataset are shown in the persistence diagram.
@@ 4 The curse of dimensionality for persistent homology
Figure 3: a – c. Persistence diagrams of a noisy circle in different ambient dimensionality and with different amount of noise. Ideally, there should be one feature (point) with high persistence, corresponding to the circle. But for high noise and dimensionality that feature vanishes into the noise cloud near the diagonal. d – f. Multidimensional scaling of Euclidean, effective resistance and diffusion distances for a noisy circle in $\mathbb{R}^{50}$. Color indicates the distance to the highlighted point.
While persistent homology is robust to small changes in point positions [17], the curse of dimensionality can still severely hurt its performance. To illustrate, we consider the same toy setting as in Figure 1: we sample points from $S^{1}\subset\mathbb{R}^{d}$, and add Gaussian noise of standard deviation $\sigma$ to each ambient coordinate. When $d=2$, higher noise does not affect the birth times but leads to lower death times (Figure 3a), because some points get distorted to the middle of the circle and the hole fills up at earlier $\tau$. When we increase the ambient dimensionality to $d=20$, higher noise leads to later birth times (Figure 3b) because in higher dimensionality distances get dominated by the noise dimensions rather than by the circular structure. Indeed, in Corollary B.5 we prove that for any two points $x_{i},x_{j}\in\mathbb{R}^{d}$ and two isotropic, multivariate normal noise vectors $\varepsilon_{1},\varepsilon_{2}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I}_{d})$ the ratio $\|\varepsilon_{1}-\varepsilon_{2}\|/\|x_{i}+\varepsilon_{1}-(x_{2}+\varepsilon_{2})\|\to 1$ in probability as $d\to\infty$. Finally, for $d=50$ both the birth and the death times increase with $\sigma$ (Figure 3c, Corollary B.7) and the ground-truth hole disappears in the cloud of spurious holes. Applying MDS to the Euclidean distances obtained with $d=50$ and $\sigma=0.25$ yields a 2D layout with almost no visible hole, because all distances have become similar (Figure 3d). See the concurrent work of Hiraoka et al. [36] for a more formal treatment.
Therefore, the failure modes of persistent homology differ between low- and high-dimensional spaces: While in low dimensions, persistent homology is susceptible to outlier points in the middle of the circle, in high dimensions, there are no points in the middle of the circle; instead, all distances become too similar, hiding the true loops. See Appendix K for more details on the effect of outliers.
@@ 5 Spectral distances are more robust
There are many possible random walks between nodes $i$ and $j$ if they both reside in the same densely connected region of the graph, while it is unlikely for a random walker to cross between sparsely connected regions. As a result, both effective resistance and diffusion distance are small between parts of the graph that are densely connected and are robust against single stray edges. This makes spectral distances on the $k$NN graph the ideal input to persistent homology for detecting the topology of data in high-dimensional spaces. Indeed, the MDS embedding of the effective resistance and of the diffusion distance of the circle in ambient $\mathbb{R}^{50}$ both clearly show the circular structure (Figure 3e,f).
@@ 6 Relation between spectral distances
We will show in Section 7 that spectral methods excel as input distances to persistent homology for high-dimensional data. But first, we will explain the relationships between them. Laplacian Eigenmaps distance and diffusion distance can be written as Euclidean distances in data representations given by appropriately scaled eigenvectors of the graph Laplacian. In this section, we derive a similar closed-form formula for effective resistance and will show that effective resistance aggregates all but the most local diffusion distances.
@@
In Appendix C we prove that the corrected effective resistance [70] can also be written in this form:
@@
It has been known [46] that the uncorrected effective resistance can be written in terms of diffusion distances as $\tilde{d}^{\textnormal{eff}}_{ij}=\sum_{t=0}^{\infty}d^{\textnormal{diff}}_{ij}(t/2)^{2}/\text{vol}(G)$, see Proposition D.1. Here, based on Proposition 6.1, we derive a similar result for the corrected effective resistance (proof in Appendix D):
@@
$d^{\textnormal{eff}}_{ij}=\sum_{t=2}^{\infty}d^{\textnormal{diff}}_{ij}(t/2)^{2}/\textnormal{vol}(G)\quad\text{ and }\quad\tilde{d}^{\textnormal{eff}}_{ij}-d^{\textnormal{eff}}_{ij}=\Big{(}d^{\textnormal{diff}}_{ij}(0)^{2}+d^{\textnormal{diff}}_{ij}(1/2)^{2}\Big{)}/\textnormal{vol}(G)$ (7)
@@
The naive effective resistance decays the eigenvectors with $1/\sqrt{\mu_{i}}$, which is much slower than diffusion distances’ $(1-\mu_{i})^{t}$ for $t\in[8,64]$. Corrected effective resistance shows intermediate behaviour (Figure 4b), because it contains some lower time diffusion distances, which decay the eigenvalues less, but not that for $t=0$, which has no decay. The correction matters little for $S^{1}\subset\mathbb{R}^{50}$ in the absence of noise, when the first eigenvalues are much smaller than the rest and dominate the embedding (Figure 4a,c) but becomes important as the noise and consequently the low eigenvalues increase (Figure 4a,d,e). As the noise increases, the decay for diffusion distances gets closer to a step function preserving only the first two non-constant eigenvectors, sufficient for the circular structure. In contrast, Laplacian Eigenmaps needs the number of components as input (Figure 4c – e). [^2] Diffusion pseudotime (DPT) [31] has a very similar expression as corrected effective resistance, using the scaling $(1-\mu_{i})/\mu_{i}$, see Appendix E. This means that DPT decays eigenvalues faster than both versions of effective resistance (Figure 4). We prove an analogous statement to Corollary 6.2 for DPT in Proposition  E.1.
@@ 7 Spectral distances find holes in high-dimensional spaces
We examined twelve distances as input to persistent homology, beyond the Euclidean distance. Full definitions are given in Appendix F. First, there are some state-of-the-art approaches for persistent homology in presence of noise and outliers. Fermat distances [23] aim to exaggerate large over small distances to incorporate the density of the data. Distance-to-measure (DTM) [2] aims for outlier robustness by combining the Euclidean distance with the distances from each point to its $k$ nearest neighbors, which are high for outliers. Similarly, the core distance used in the HDBSCAN algorithm [9, 19] raises each Euclidean distance at least to the distance between incident points and their $k$-th nearest neighbors. We evaluate these methods here with respect to Gaussian noise in high-dimensional ambient space, a different noise model than the one for which these method were designed. Second, we consider some non-spectral graph distances. The geodesic distance on the $k$NN graph was popularized by Isomap [65] and used for persistent homology by Naitzat et al. [52]. Following Gardner et al. [26] we used distances based on UMAP affinities, and also experimented with $t$-SNE affinities. Third, we computed $t$-SNE and UMAP embeddings and used distances in the 2D embedding space. Finally, we explored methods using the spectral decomposition of the $k$NN graph Laplacian, see Section 6: effective resistance, diffusion distance, and the distance in Laplacian Eigenmaps’ embedding space.
@@
All methods come with hyperparameters. We report the results for the best hyperparameter setting on each dataset (Appendix H) but found spectral methods to be robust to these choices (Appendix I).
Performance metric
@@
The output of persistent homology is a persistence diagram showing birth and death times for all detected holes. It may be difficult to decide whether this procedure has actually ‘detected’ a hole in the data. Ideally, for a dataset with $m$ ground-truth holes, the persistence diagram should have $m$ points with high persistence while all other points should have low persistence and lie close to the diagonal. Therefore, for $m$ ground-truth features, our hole detection score $s_{m}\in[0,1]$ is the relative gap between the persistences $p_{m}$ and $p_{m+1}$ of the $m$-th and $(m+1)$-th most persistent features: $s_{m}=(p_{m}-p_{m+1})/p_{m}$. This corresponds to the visual gap between them in the persistence diagram (Figure 2b). Rieck and Leitte [57] used a similar quantity to find important features.
@@
In addition, we set $s_{m}=0$ if all features in the persistence diagram have very low death-to-birth ratios $\tau_{d}/\tau_{b}<1.25$. This handles situations with very few detected holes that die very quickly after being born, which otherwise can have spuriously high $s_{m}$ values. This was done everywhere apart from the qualitative Figures 1, 8 and in Figure S21. We call this heuristic thresholding.

### v3

@@ 1 Introduction
Algebraic topology can describe the shape of a continuous manifold. In particular, it can detect if a manifold has holes, using its so-called homology groups [39]. For example, a cup has a single one-dimensional hole, or loop (its handle), whereas a football has a single two-dimensional hole, or void (its hollow interior). These global topological properties are often helpful for understanding an object’s overall structure. However, real-world datasets are typically given as point clouds, a discrete set of points sampled from an underlying manifold. In this setting, true homologies are trivial, as there is one connected component per point and no holes whatsoever; instead, persistent homology can be used to find holes in point clouds and to assign an importance score called persistence to each [25, 90]. Holes with high persistence are indicative of holes in the underlying manifold. Persistence homology has been successfully applied in machine learning pipelines, for instance for gait recognition [48], instance segmentation [44], and protein binding [84], as well as for neural network analysis [67].
Persistent homology works well for low-dimensional data [78] but we find that it has difficulties in high dimensionality. If data points are sampled from a low-dimensional manifold embedded in a high-dimensional ambient space (manifold hypothesis), then the measurement noise typically affects all ambient dimensions. In this setting, traditional persistent homology is not robust against even low levels of noise. On a dataset as simple as a circle in $\mathbb{R}^{50}$, persistent homology based on the Euclidean distance between noisy points can fail to identify the correct loop as a clear outlier in the persistence diagram (Figure 1). The aim of our work is to find alternatives to traditional persistent homology that can robustly detect the correct topology despite high-dimensional noise.
@@
Our contributions are:
@@
Figure 1: a. 2D PCA of a noisy circle ($\sigma=0.25$, radius 1) in $\mathbb{R}^{50}$. Overlaid are representative cycles of the most persistent loops. b. Persistence diagrams using Euclidean distance and the effective resistance. c. Loop detection scores of persistent homology using effective resistance and Euclidean distance. d, e. UMAP and $t$-SNE embeddings of the same data, showing the loop structure in 2D.
@@ 2 Related work
Persistent homology has long been known to be sensitive to outliers [15] and several extensions have been proposed to make it more robust. One recurring idea is to replace the Euclidean distance with a different distance matrix, before running persistent homology. Bendich et al. [6] suggested to use diffusion distances [20], but their empirical validation was limited to a single dataset in 2D. Anai et al. [2] suggested to use the distance-to-measure (DTM) [15] and Fernández et al. [26] proposed to use Fermat distances [34]. Vishwanath et al. [80] introduced persistent homology based on robust kernel density estimation, an approach that itself becomes challenging in high dimensionality. All of these works focused on low-dimensional datasets ($<$10D, mostly 2D or 3D), while our work specifically addresses the challenges of persistent homology in high dimensionality.
The concurrent work of Hiraoka et al. [42] is the most relevant related work. Their treatment of the curse of dimensionality of persistent homology is mostly theoretical, while ours has an empirical focus. The two works are thus complementary to each other. Hiraoka et al.’s theoretical description of the curse of dimensionality is similar to ours (Appendix B) in that it analyses how distance concentration with high-dimensional noise impairs persistent homology, but is more general. Practically, Hiraoka et al. propose normalized PCA to mitigate the curse of dimensionality. However, this approach assumes the true dimensionality of the data to be known, which is not realistic in real-world applications, and performs worse than our suggestions (Appendix C).
Below, we recommend using effective resistance and diffusion distances for persistent homology in high-dimensional spaces. Both of these distances, as well as the shortest path distance, have been used in combination with persistent homology to analyze the topology of graph data [62, 37, 1, 77, 11, 55, 23]. Shortest paths on the $k$NN graph were also used by Naitzat et al. [60] and Fernández et al. [26]. Motivated by the performance of UMAP [53] for dimensionality reduction,  Gardner et al. [31] and Hermansen et al. [40] used UMAP affinities to define distances for persistent homology.
Effective resistance is a well-established graph distance [24, 29]. A correction, more appropriate for large graphs, was suggested by von Luxburg et al. [81, 83]. When speaking of effective resistance, we mean this corrected version, if not otherwise stated. It has not yet been combined with persistent homology. Conceptually similar diffusion distances [20] have been used in single-cell RNA-sequencing data analysis, for dimensionality reduction [56], trajectory inference [35], feature extraction [16], and hierarchical clustering, similar to 0D persistent homology [9, 46].
@@
Figure 2: a. Persistent homology applied to a noisy circle ($n=10$) in 2D tracks appearing and disappearing holes as balls grow around each datapoint. Dotted lines show the graph edges that lead to the birth / death of two loops (Section 3). b. The corresponding persistence diagram with two detected 1D holes (loops). Our hole detection score measures the gap in persistence between the first and the second detected holes (Section 7).
@@ 3 Background: persistent homology
Persistent homology consists of a set of holes for each dimension. We limit ourselves to loops and voids. Each hole has associated birth and death times $(\tau_{b},\tau_{d})$, i.e., the first and last filtration value $\tau$ at which that hole exists. Their difference $p=\tau_{d}-\tau_{b}$ is called the persistence of the hole and quantifies its prominence. The birth and death times can be visualized as a scatter plot (Figure 2b), known as the persistence diagram. Points far from the diagonal have high persistence. This process is illustrated in Figure 2 for a noisy sample of $n=10$ points from a circle $S^{1}\subset\mathbb{R}^{2}$. At $\tau_{1}$, a small spurious loop is formed thanks to the inclusion of the dotted edge, but it dies soon afterwards. The ground-truth loop is formed at $\tau_{2}$ and dies at $\tau_{3}$, once the hole is completely filled in by triangles. Both loops (one-dimensional holes) found in this dataset are shown in the persistence diagram.
@@ 4 The curse of dimensionality for persistent homology
Figure 3: a – c. Persistence diagrams of a noisy circle in different ambient dimensionality and with different amount of noise. Ideally, there should be one feature (point) with high persistence, corresponding to the circle. But for high noise and dimensionality that feature vanishes into the noise cloud near the diagonal. d – f. Multidimensional scaling of Euclidean, effective resistance, and diffusion distances for a noisy circle in $\mathbb{R}^{50}$. Color indicates the distance to the highlighted point.
While persistent homology is robust to small changes in point positions [19], the curse of dimensionality can still severely hurt its performance. To illustrate, we consider the same toy setting as in Figure 1: we sample points from $S^{1}\subset\mathbb{R}^{d}$, and add Gaussian noise of standard deviation $\sigma$ to each ambient coordinate. When $d=2$, higher noise does not affect the birth times but leads to lower death times (Figure 3a), because some points get distorted to the middle of the circle and the hole fills up at earlier $\tau$. When we increase the ambient dimensionality to $d=20$, higher noise leads to later birth times (Figure 3b) because in higher dimensionality distances get dominated by the noise dimensions rather than by the circular structure. Indeed, in Corollary B.5 we prove that for any two points $x_{i},x_{j}\in\mathbb{R}^{d}$ and two isotropic, multivariate normal noise vectors $\varepsilon_{1},\varepsilon_{2}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I}_{d})$ the ratio $\|\varepsilon_{1}-\varepsilon_{2}\|/\|x_{i}+\varepsilon_{1}-(x_{2}+\varepsilon_{2})\|\to 1$ in probability as $d\to\infty$. Finally, for $d=50$ both the birth and the death times increase with $\sigma$ (Figure 3c, Corollary B.7) and the ground-truth hole disappears in the cloud of spurious holes. Applying MDS to the Euclidean distances obtained with $d=50$ and $\sigma=0.25$ yields a 2D layout with almost no visible hole, because all distances have become similar (Figure 3d). See the concurrent work of Hiraoka et al. [42] for a more detailed treatment.
Therefore, the failure modes of persistent homology differ between low- and high-dimensional spaces. While in low dimensions, persistent homology is susceptible to outlier points in the middle of the circle, in high dimensions, there are no points in the middle of the circle; instead, all distances become too similar, hiding the true loops. See Appendix N for more details on the effect of outliers.
@@ 5 Spectral distances are more robust
There are many possible random walks between nodes $i$ and $j$ if they both reside in the same densely connected region of the graph, while it is unlikely for a random walker to cross between sparsely connected regions. As a result, both effective resistance and diffusion distance are small between parts of the graph that are densely connected and are robust against single stray edges (Figure 4). This makes spectral distances on the $k$NN graph the ideal input to persistent homology for detecting the topology of data in high-dimensional spaces. Indeed, the MDS embedding of the effective resistance and of the diffusion distance of the circle in ambient $\mathbb{R}^{50}$ both clearly show the circular structure (Figure 3e,f).
[…]
@@ 6 Relation between spectral distances
We show in Section 7 that spectral methods excel as input distances to persistent homology for high-dimensional data. But first, we explain the relationships between them. Laplacian Eigenmaps distance and diffusion distance can be written as Euclidean distances in data representations given by appropriately scaled eigenvectors of the graph Laplacian. In this section, we derive a similar closed-form formula for effective resistance and show that effective resistance aggregates all but the most local diffusion distances.
@@
In Appendix F we prove that the corrected effective resistance [81] can also be written in this form:
@@
It has been known [54] that the uncorrected effective resistance can be written in terms of diffusion distances as $\tilde{d}^{\textnormal{eff}}_{ij}=\sum_{t=0}^{\infty}d^{\textnormal{diff}}_{ij}(t/2)^{2}/\text{vol}(G)$, see Proposition G.1. Here, based on Proposition 6.1, we derive a similar result for the corrected effective resistance (proof in Appendix G):
@@
$d^{\textnormal{eff}}_{ij}=\sum_{t=2}^{\infty}d^{\textnormal{diff}}_{ij}(t/2)^{2}/\textnormal{vol}(G)\quad\text{ and hence }\quad\tilde{d}^{\textnormal{eff}}_{ij}-d^{\textnormal{eff}}_{ij}=\Big{(}d^{\textnormal{diff}}_{ij}(0)^{2}+d^{\textnormal{diff}}_{ij}(1/2)^{2}\Big{)}/\textnormal{vol}(G).$ (7)
@@
The naive effective resistance decays the eigenvectors with $1/\sqrt{\mu_{i}}$, which is much slower than diffusion distances’ $(1-\mu_{i})^{t}$ for $t\in[8,64]$. Corrected effective resistance shows intermediate behavior (Figure 5b). When represented as a sum over diffusion distances, it contains all diffusion distances with $t\geq 1$, making it decay slower than diffusion distances with $t=8$ or $64$, but does not contain the non-decaying $t=0$ term, so it decays faster than its naive version. The correction matters little for $S^{1}\subset\mathbb{R}^{50}$ in the absence of noise, when the first eigenvalues are much smaller than the rest and dominate the embedding (Figure 5a,c) but becomes important as the noise and consequently the low eigenvalues increase (Figure 5a,d,e). As the noise increases, the decay for diffusion distances gets closer to a step function preserving only the first two non-constant eigenvectors, sufficient for the circular structure. In contrast, Laplacian Eigenmaps needs the number of components as input (Figure 5c – e). [^1] Diffusion pseudotime (DPT) [36] has a very similar expression as corrected effective resistance, using the scaling $(1-\mu_{i})/\mu_{i}$, see Appendix H. This means that DPT decays eigenvalues faster than both versions of effective resistance (Figure 5). We prove an analogous statement to Corollary 6.2 for DPT in Proposition  H.1.
@@ 7 Spectral distances find holes in high-dimensional spaces
We examined twelve distances as input to persistent homology, beyond the Euclidean distance. Full definitions are given in Appendix I. First, there are some state-of-the-art approaches for persistent homology in the presence of noise and outliers. Fermat distances [26] aim to exaggerate large over small distances to incorporate the density of the data. Distance-to-measure (DTM) [2] aims for outlier robustness by combining the Euclidean distance with the distances from each point to its $k$ nearest neighbors, which are high for outliers. Similarly, the core distance used in the HDBSCAN algorithm [10, 21] raises each Euclidean distance at least to the distance between incident points and their $k$-th nearest neighbors. We evaluate these methods here with respect to Gaussian noise in high-dimensional ambient space, a different noise model than the one for which these methods were designed. Second, we consider some non-spectral graph distances. The geodesic distance on the $k$NN graph was popularized by Isomap [76] and used for persistent homology by Naitzat et al. [60]. Following Gardner et al. [31] we used distances based on UMAP affinities, and also experimented with $t$-SNE affinities. Third, we computed $t$-SNE and UMAP embeddings and used distances in the 2D embedding space. Finally, we explored methods using the spectral decomposition of the $k$NN graph Laplacian, see Section 6: effective resistance, diffusion distance, and the distance in Laplacian Eigenmaps’ embedding space.
@@
All methods come with hyperparameters. We report the results for the best hyperparameter setting on each dataset (Appendix K) but found spectral methods to be robust to these choices (Appendix L).
Performance score
@@
The output of persistent homology is a persistence diagram showing birth and death times for all detected holes. It may be difficult to decide whether this procedure has actually detected a hole in the data, or not. Ideally, for a dataset with $m$ ground-truth holes, the persistence diagram should have $m$ points with high persistence while all other points should have low persistence and lie close to the diagonal. Therefore, for $m$ ground-truth features, our hole detection score $s_{m}\in[0,1]$ is the relative gap between the persistences $p_{m}$ and $p_{m+1}$ of the $m$-th and $(m+1)$-th most persistent features: $s_{m}=(p_{m}-p_{m+1})/p_{m}$. This corresponds to the visual gap between them in the persistence diagram (Figure 2b). Rieck and Leitte [66] as well as Smith and Kurlin [74] used similar quantities to find important features. We prove a continuity property of $s_{m}$ in Appendix D and consider alternative scores in Appendix E.
@@
In addition, we set $s_{m}=0$ if all features in the persistence diagram have very low death-to-birth ratios $\tau_{d}/\tau_{b}<1.25$. This handles situations with very few detected holes that die very quickly after being born, which otherwise can have spuriously high $s_{m}$ values. This was done everywhere apart from the qualitative Figures 1, 9 and in Figure S25. We call this heuristic thresholding.